A scientist wants to make 6 milliliters of a 30% alcohol solution. The solution is to be made from a combination of a 20% alcohol solution and a 50% alcohol solution.
a. Write a system of two equations that model this situation (Define variables)
b. How many milliliters of each solution must be combined to get 6 milliliters of 30% alcohol solution?

Answer by richwmiller(17219)

We need a total of 6 mL at 30%
We want to know how many mL at 20% to mix with 50%
a+b=6,
0.2*a+0.5*b=0.3*6
a=6-b
0.2*(6-b)+0.5*b=1.8
1.2-0.2b+0.5*b=1.8
0.3*b=0.6
b=2
a=6-b
a=4 mL at 20%
b=2 mL at 50%
check
0.2*4+0.5*2=0.3*6
0.8+1=1.8
1.8=1.8
ok
codemx1